Numerical Analysis I, Fall 2009 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 11

Assigned Nov 27, 2009. Due Dec 04, 2009.

1. Section 5.3: Problems 2(d,e,f), 9, 11.

For problem 11 (b), just write down the equations for the unknowns and take the answer for granted. Need not verify it.

2. Section 5.4: Problems 6, 8, 10, 11.

For problem 10: These formula, also known as one-sided difference, can be used to find approximate derivatives at the left end point x_{0}, given f(x_{0}), f(x_{1}), ⋯, f(x_{0}). Also write down similar formula at x_{n} by symmetry. Need not re-derive the formula there again.

For problem 11, read section 5.4.3 first.

3. Challenge of the week: later.